# Diffusion MRI with Machine Learning

### v1

@@ II-F Segmentation
Tissue segmentation in the dMRI space is needed for important applications such as anatomically constrained tractography and structural connectivity analysis [133, 257, 138]. Given the lower spatial resolution and tissue contrast of dMRI, this segmentation is usually obtained via segmenting anatomical MRI data followed by registration to the dMRI space. This registration can introduce significant error because the contrast and image distortions between the modalities can be very different. Therefore, direct tissue segmentation based on the dMRI signal is highly desirable. Machine learning methods, and in particular deep learning techniques, have shown great promise here as well.
@@
Prior to the development of deep learning-based segmentation methods, several works applied classical machine learning methods for tissue segmentation in dMRI. Fuzzy c-means clustering with spatial constraints [258] and sparse representation of the dMRI signal in dictionaries [259] were successfully used to segment the brain tissue in the dMRI space. Ciritsis et al. determined the optimal set of dMRI measurements and DTI features for tissue classification with a support vector machines (SVM) method [260]. Schnell et al., on the other hand, computed rotation-invariant features from HARDI acquisitions as input to an SVM to segment the brain tissue into six classes, which included white matter with parallel (single) fibers and white matter with crossing fibers [261].
@@
More recently, Golkov et al. [26] extended their deep learning micro-structure estimation technique to develop a segmentation method that mapped the dMRI signal in each voxel to a probability map for WM/GM/CSF and an MS lesion class. Their simple and flexible approach allowed them to include an extra FLAIR channel, which could be especially useful for MS lesion segmentation. Their model accurately segmented the brain tissue, and detected MS lesion with AUC in the range 0.88-0.94. The same authors also proposed a variational auto-encoder model for novelty detection based on q-space dMRI data and showed that it could accurately segment MS lesions with an AUC of approximately 0.90 [262]. Another study used DTI and DKI features as input to a multi-view CNN to achieve accurate segmentation of WM, GM, and CSF [263]. The model was trained on HCP-quality data, where the anatomical and dMRI data could be registered precisely. The trained model achieved high segmentation accuracy on clinical dMRI acquisitions. Similar results were reported with DKI features and a U-Net model [264]. For isointense stage of brain development where white matter and gray matter have similar intensity in T1 and T2 images (approximately between 6 and 8 months), one study used an additional FA channel to accurately segment the brain tissue with an FCN [265]. Segmentation of the intracranial volume also know as brain masking or skull stripping, which is typically performed as a pre-processing step, has also been addressed with machine learning methods [266, 267]. It has been shown that deep learning methods can achieve better accuracy than other techniques [268, 269, 267].
II-G Quality enhancement and interpolation
Increasing the spatial resolution or the q-space sampling density would mean increasing the scan time. As an alternative, recent studies have proposed to use computational methods. Some of the proposed techniques have been inspired by deep learning-based super-resolution techniques. However, often these standard deep learning techniques have to be significantly modified to handle the dMRI data. We provide a summary of some of these methods and their findings below.
  • Estimating unobserved q-space shells from observed ones has been reported by several works [270, 271, 92]. Murray et al. used neural networks to predict multi-shell HARDI data from single-shell data and used the predicted data to estimate the NODDI model [270]. They achieved satisfactory results on healthy brains as well as brain of MS patients. In a similar work, Chen et al. [92] successfully trained an MLP to predict six-shell data from two-shell data.
  • Several works have used neural networks to enhance the angular resolution of q-space data [92, 272, 273, 274]. Yin et al. used a method based on sparse representation of the dMRI signal and a neural network built on 1D convolutions to map low angular resolution data to high angular resolution data. Evaluations showed that this method improved the FOD reconstruction accuracy for complex fiber configurations [272]. Lyon et al. proposed a parametric continuous convolutional network for angular super-resolution of dMRI data and showed that this approach improved the results of complex computations such as fixel-based analysis and NODDI model fitting [274]. Another work has proposed a flexible model, based on a recurrent convolutional neural network (RCNN), that can predict the unobserved data for arbitrary target b-vectors [275]. To achieve this flexibility, the model follows an encoder-decoder design where the encoder uses the gradient table information of the measured data as additional input. The decoder, then, takes the latent representation generated by the encoder and the gradient table of the target (unobserved) data to make the prediction. It has been shown that this model can accurately upsample the dMRI data by a factor of 8, and that it is more accurate than predictions based on spherical harmonic interpolation and a 1D variant of the model architecture. Experiments have also shown that at higher sub-sampling rates spatial information can substantially improve the accuracy of angular interpolation [275]. Similarly, Ren et al. have developed a method for predicting diffusion-weighted images for arbitrary gradient directions [276]. The input to their model includes only T1, T2, and b0 (non-diffusion-weighted) images. Methodologically, their model is a generative adversarial network (GAN, [277]) where both the generator and the discriminator are conditioned on the gradient direction and strength. Experiments are performed on training and test data from HCP. The results show that the synthesized data can be used for estimating DKI, NODDI, and FOD.
  • Different from the methods above, a number of works have attempted to increase the spatial resolution of dMRI data in the voxel space [278, 279]. Still others have proposed to learn to enhance the quality/resolution of the target parameter maps. Zeng et al. developed a method to enhance the resolution of FOD maps [280]. Their method used the FOD estimated from low angular resolution single-shell data to predict those computed from multi-shell HARDI data. Experiments on clinical quality measurements showed that this method significantly improved the tractography and structural connectivity assessment. A similar study has been reported by Lucena et al., [281], where the authors trained CNNs to map the FODs computed from single-shell data to those computed from multi-shell data. The method worked well and displayed satisfactory generalizability to different scanners and data acquisition protocols. Ye et al. [93], on the other hand, proposed to combine a learning-based parameter mapping method similar to [62] with a super-resolution method similar to [282] in a joint framework to estimate high-resolution NODDI maps from dMRI measurements that had lower resolution in both voxel space and q-space. It is also possible to jointly improve the spatial resolution and q-space resolution, as proposed by Ye at al. in their super-resolved q-space deep learning (SR-q-DL) framework [93, 283, 29]. In this framework, patches from the source low-resolution and the target high-resolution dMRI volumes are represented in dictionaries and a neural network is trained to map the representation of the source signal to that of the target signal. This approach has been shown to lead to significant improvements in the accuracy of NODDI estimation from low-quality dMRI data [93, 283]. In a more recent work, Spears and Fletcher [284] have argued that prior voxel-space super-resolution methods such as [29] are not ideal for tractography, where a continuous FOD field is desired. Instead, they have proposed a method based on continuous and differentiable signal representations with neural networks. They show that this method achieves state of the art FOD estimation and tractography on HCP data.
  • In a series of works, Alexander et al. have developed and promoted the notion of image quality transfer (IQT) that uses machine learning methods such as random forests and neural networks to improve the quality of dMRI parameter maps [285, 286, 287, 288]. The goal of IQT is to use high-quality data such as the HCP data to improve the quality of regular dMRI data that are typically obtained in clinical applications. The method works by learning a regression function to map the low-quality dMRI patches to high-quality patches. This idea has been successfully applied to increase the spatial resolution of diffusion tensor images and to estimate NODDI parameters from low-quality single-shell measurements. Furthermore, it has been shown that on low-quality data, IQT enables tractography-based reconstruction of thin fiber tracts that can typically only be reconstructed using specialized data.
  • Another related and highly important application is that of imputing missing data. This is of particular importance since some dMRI measurements often have to be discarded due to strong noise, artifacts, or excessive motion, especially in young or noncooperative subjects. To address this problem, Hong et al. developed a method based on graph CNNs [31]. They represented the relation between the measurements in the voxel space and q-space using a graph, and then trained a residual graph CNN to learn this relation. They trained the model using adversarial techniques and successfully applied it on data from healthy infants between 0 and 12 months old. A rather similar technique was proposed to enable reduction of scan time in [289]. Specifically, the authors proposed to measure only a subset of slices in each dMRI volume and learn to predict the unmeasured slices with a deep learning model based on graph CNNs. To achieve superior reconstruction quality, a generative adversarial approach was used to train the model. This method was able to reconstruct accurate maps of generalized fractional anisotropy (GFA) and FOD from five-fold-reduced slice acquisitions.
II-H Data harmonization
Data harmonization is crucial for reliable analysis of inter-scanner and inter-site dMRI data. It has become an increasingly more relevant problem as larger and geographically/demographically more diverse datasets are used in multi-center neuroimaging studies. Statistical methods such as ComBat [290] and RISH [291] have achieved considerable success in addressing this problem. However, they have important limitations. For example, they depend on data from matched subjects in the reference and target datasets, and their performance deteriorates when the differences between reference and target datasets are nonlinear [292]. Recently, various machine learning-based methods have been proposed, but they still have not achieved the popularity of methods such as ComBat and RISH. A recent study compared a range of interpolation methods, statistical regression techniques, and deep learning methods for cross-scanner and cross-protocol multi-shell data harmonization [293]. A regression method based on [294] performed better than all deep learning methods. However, some of the deep learning methods were among the best performing techniques. The authors of [293] hypothesized that the performance of the deep learning methods may significantly improve with larger training datasets than the 10 training subjects used in that work. Tax et al. evaluated five different learning-based harmonization methods, including four neural networks and a dictionary-based method, on data from different scanners and with different maximum gradient strengths [295]. Their results showed that overall these learning-based methods were successful in harmonizing cross-scanner and cross-protocol data, although no comparison with the state of the art non-learning methods was reported. Their analysis also showed that the learning-based methods were more effective on isotropic measures such as MD than on anisotropic measures such as FA. They attribute this behavior to the possibility that, because the spatial variations in isotropic measures are less abrupt, imperfect spatial alignments may be less harmful when applying machine learning methods on registered pairs of data.
Table VIII lists some of the proposed machine learning methods. In general, these methods offer higher flexibility in modeling the heterogeneity in the data and in handling data from unpaired subjects. For example, one study developed a method based on disentangled latent space representation of the data that allowed for separating the effects of anatomy and acquisition [296]. Similarly, variational auto-encoders have been proposed to harmonize dMRI data by learning a latent representation that is invariant to site/protocol-specific factors [297]. Focusing on the specific task of FOD estimation, another study developed a method, named null space deep network, that could seamlessly integrate a small dMRI dataset with histology ground truth and a large scan-rescan dMRI dataset without histology ground truth [298]. The resulting method showed superior FOD estimation accuracy on data from scanners that had not been included in the training set. A similar method was applied to harmonize structural connectivity metrics on a two-center dataset [299]. The authors found that optimal harmonization was achieved when the deep learning method was applied to harmonize the dMRI data at a voxel level followed by applying ComBat [290] to harmonize the structural connectivity metrics. Another work from the same research team, again focusing on FOD estimation, proposed an additional loss term to encourage consistency of FOD estimations computed from scan-rescan data of the same subject [122]. Evaluations showed that the new loss function improved the generalizability of FOD estimation as well as downstream tasks such as structural connectivity assessment. Blumberg et al. have proposed a multi-task learning strategy, where a neural network model is trained to combine the information learned by an ensemble of different neural network models [300]. The individual neural nets in the ensemble may have been trained on separate datasets, possibly for completely different tasks. The added neural network model uses the high-level features learned by the ensemble to perform data harmonization. The authors argue that this approach can be especially effective in scenarios where one or some of the datasets are very small.
TABLE VIII: A listing of some of the machine learning-based methods for data harmonization in dMRI.
 Reference
Method
Main results
 [296]
semi-supervised contrastive learning using CNNs
This method uses paired data to learn subject-specific and acquisition-specific representations. A decoder is trained to map the subject-specific representations to the target contrast. The method is shown to be superior to a method based on CycleGAN as well as interpolation based on SHORE. It can handle heterogeneous data from multiple sites, different acquisitions protocols, and demographics.
[298, 301]
null space learning with an MLP architecture
The model is trained using paired dMRI scans of human subjects as well as dMRI-histology data from squirrel monkeys. The loss function is designed to encourage accurate FOD estimation and consistent scan-rescan GFA. Experiments show that this method has high accuracy in FOD estimation and high reproducibility on test data from unseen scanners. Other experiments show that this approach results in higher FOD estimation accuracy across 1.5T and 3T scanners.
[297]
variational auto-encoder
The method is based on learning a latent representation that is invariant to scanner, scanning protocol, or similar confounders. The latent representation can then be used to reconstruct the image content that is stripped of those factors. The method can be trained without paired scans. Training involves an adversarial loss that attempts to predict the source (scanner, etc.) of the source image. Results show that, compared with a non-deep learning method [302], this new method achieves superior results in terms of several parameters including FA, MD, and fiber orientation.
[303]
Residual CNN
The method is based on predicting the spherical harmonics representation of the dMRI signal in the target domain from that of source domain. It works on 3D patches of size 3. A final non-learned projection in the spherical harmonics space is needed to make sure fiber orientations are not changed. This method requires that the same subjects be scanned in the source and target scanner/acquisition protocol. Evaluations show that this method can achieve effective harmonization of the dMRI signal, FA, and MD.
[304]
hierarchical CNNs
The aim of this work is on harmonization of DKI measures. The method requires a set of subjects to be scanned with both target and source scanners/acquisition protocols. The method computes the DKI metrics using the data in the target domain. The network is trained to map the dMRI data in the source domain directly to the DKI measures in the target domain, after nonlinear intra-subject registration. Experiments show that this method leads to 51-66% reduction in inter-scanner variation of DKI measures.
[122, 305]
CNN with a scan-rescan consistency loss
This method is proposed and evaluated specifically for FOD estimation, although it seem to be directly applicable for estimating any other parameter. The consistency loss penalizes the model for divergent FOD predictions from different scans of the same subject. The method outperforms standard methods on external test data. It achieves improved FOD estimation accuracy and intra-scanner consistency.
[300]
ensemble of different neural networks
The authors present their method as a multi-task learning approach. Essentially, a set of neural networks, which can have different architectures and trained for different prediction tasks are combines via training a set of additional neural networks that utilize the features learned by these networks to predict the parameter of interest. Compared with several state of the art deep learning methods, this method resulted in superior dMRI signal prediction.
[306]
adaptive dictionary learning
A dictionary is learned on set of target datasets/scanner(s). It is assumed that representing a test scanners data in this dictionary automatically harmonizes the data towards the target dictionary by suppressing the features that are specific to the test data. The method does not require paired subjects in the source and target datasets. Successful harmonizations are reported in terms of FA, MD, and rotationally invariant spherical harmonics representations.
II-I Data preprocessing
Diffusion-weighted MRI measurements suffer from various artifacts and imperfections that can influence the results of downstream analysis steps and the conclusions drawn from the data. Therefore, data preprocessing is an essential first step for any dMRI analysis [4]. Although the main focus of this survey is on post-processing analysis, in this section we briefly explain some of the recent progress in using machine learning methods to improve pre-processing of dMRI data.
II-I1 Denoising
Measurement noise in dMRI has a spatially-varying distribution and depends on factors such as acquisition and reconstruction methods. It is typically approximated using Rician or non–central chi–square distributions [307, 308]. The noise is stronger at high b-values that are essential for estimating fiber orientation and multi-compartment models and it is of high practical importance because it can influence the analysis results in subtle but significant and complex ways [309, 8, 9]. Denoising methods based on the principal components analysis and non-local means have been popular in dMRI [310, 311]. These methods were often adapted to the dMRI setting in innovative ways. For example St-Jean et al. [312] combined the non-local matching with dictionary learning. In order to adapt the non-local means method to curved non-Euclidean domain of dMRI signals, Chen et al. [313] developed a method based on convolutions on graphs that enabled computing multi-scale features from dMRI data.
In recent years, several neural network-based image denoising algorithms ave been proposed by the computer vision society. These methods, such as Noise2Noise [314], Noise2Void [315], and Noise2Self [316] have shown to match or surpass the performance of classical denoising algorithms and offer several advantages that are especially useful for medical imaging applications including dMRI: (1) They do not need to model the image prior; (2) They do not need to know the noise distribution; (3) They do not require clean sample images or paired noisy-clean images for training. The ideas presented in these methods have been adopted to develop dMRI denoising techniques in a few recent studies [317, 318]. Fadnavis et al. proposed Patch2Self, which is based on learning locally-linear relations between different dMRI volumes [317]. Patch2Self predicts the signal in one of the q-space elements (i.e., channels in a 4D representation of the dMRI data) based on the other elements. In order to exploit spatial correlations, the model uses 3D spatial patches as input and target. Effectiveness of Patch2Self has been demonstrated in a study on spinal cord [319], where it has been shown that Patch2Self improves FA repeatability and MS lesion conspicuity. SDnDTI [318] is another self-supervised method that uses a CNN model. It is specially tailored to single-shell data for DTI estimation. Given a set of N dMRI volumes (each acquired with a different gradient direction but the same gradient strength), SDnDTI selects subsets of six volumes and denoises them. This is done by training the CNN to take these six volumes as input and predict a low-noise target that is computed from all N volumes. This is repeated until all N volumes are denoised. Experiments show that SDnDTI is superior to standard methods such as those based on PCA and non-local means.
II-I2 Artifact correction
Diffusion MRI measurements can also suffer from a range of artifacts such as those caused by off-resonance fields, subject motion, and Gibbs ringing [7, 320]. In the past two decades, much research has been devoted to computational methods for retrospective correction of these artifacts. A recent review of these methods has been written by [4]. Machine learning methods have also been proposed to address these artifacts.
Much of the distortion correction algorithms for dMRI, such as those targeting susceptibility-induced distortions, are based on image registration. Image registration for dMRI data is especially challenging because of several factors: (1) Low SNR, especially for higher b-values; (2) Highly different appearance of dMRI volumes of the same brain acquired with different gradient strengths and directions, which is again exacerbated for higher b values, making it difficult to design registration methods; (3) Difficulty of finding a reference image for registration since all dMRI volumes are affected by distortions; (4) Complex interaction of various artifacts such as signal loss that can further challenge the registration process. Standard methods for distortion correction, such as topup [321], use iterative optimization to compute the displacement field. Deep learning-based registration methods have the potential to address these challenges because they can be trained in an unsupervised manner on data with various imperfections.
Unsupervised [322, 323, 324], semi-supervised [325], and supervised [30] deep learning-based registration methods have been applied on reversed phase-encoding dMRI volumes to estimate the displacement field. One study [325] combined unsupervised training with supervised training using the deformation fields computed by topup as the target. Their experiments showed that their proposed semi-supervised method could lead to better results than purely supervised or purely unsupervised techniques. Most prior works have shown that the deep learning methods can match the accuracy of the conventional methods such as topup while being orders of magnitude faster [322, 324, 325, 323, 326]. One study, used distortion-free volumes obtained with point-spread-function encoded EPI as registration target and achieved more accurate distortion correction than topup [30]. The input to their network, which was a 2D FCN, consisted of distorted dMRI volumes as well as an anatomical T2 image, and the network output was undistorted dMRI volumes. Deep learning methods may also have good generalizability. Some works have shown that the deep learning models trained on normal brains work well on abnormal brains [30, 323] and also on different (left-right) phase encoding directions [323].
While most deep learning-based methods have used the dMRI volumes to compute the displacements for distortion correction, Qiao et al. propose to use FODs instead [327, 328]. They argue that the information contained in dMRI data volumes is not sufficient to resolve detailed white matter structures that are essential for distortion correction in areas such as the brainstem [327]. Instead, they compute FOD maps from dMRI volumes acquired with opposite phase encodings and train a U-Net to compute the displacement fields. They show that this method achieves better results than topup and two deep learning methods in the brainstem and white matter. Another interesting work is that of Schilling et al. [326, 329], which enables susceptibility-induced distortion correction when scans wit reversed phase-encodings is not available. They have developed a deep learning models (an FCN or a GAN) to generate a undistorted b0 (non-diffusion-weighted) volume given (i) a distorted b0 volume and (ii) a anatomical (T1) image. The undistorted and distorted image can then be used to compute the displacement field with topup.
Deep learning methods have been developed to address other types of artifacts in dMRI. Ayub et al. have proposed a method based on variational auto-encoders for in-painting cropped dMRI volumes [330]. The model, trained on artificially-cropped volumes and tested on real-world cropped data, was able to correctly inpaint the cropped data. Deep learning methods have also been successfully applied to remove Gibbs artifacts in dMRI data. One study trained a CNN on T2 images to estimate Gibbs ringing artifacts [331]. The trained model corrected the Gibbs artifacts in dMRI volumes as well. Another work trained a CNN on natural images with simulated artifacts and showed that the trained model effectively suppressed Gibbs artifacts in dMRI data and in the computed parameter maps [332].
II-I3 Automatic quality control
Quality control in dMRI is challenging due to the diversity of artifacts and the large data size. A number of deep learning methods have been proposed for automatic quality control in large dMRI studies. Ahmad et al. [333] trained a CNN to classify dMRI volumes as normal or containing artifacts. The artifacts considered in their study included motion-induced signal dropout, inter-slice instability, ghosting, chemical shift and susceptibility-induced distortions. Their method showed an accuracy of 92% on a pool of seven datasets with different acquisition protocols. Another work used a mix of real and simulated data to develop a quality control method, based on CNNs, for detecting intra-volume motion and signal dropout [334]. To detect motion-corrupted volumes in neonatal dMRI volumes, Kelly et al. trained a random forest on the outputs of an ensemble of CNNs and achieved classification accuracies of well above 90% [335]. Samani et al. also trained a CNN to detect a range of artifacts such as motion artifacts, ghosting, susceptibility-induced artifacts, and chemical shift [336]. They reported a detection accuracy of 98% on multi-scanner dMRI data. Both [334] and [336] used transfer learning to train their networks.
III Discussion
@@
In this section, we present our thoughts regarding the main aspects of the applications of machine learning for dMRI analysis. We discuss some of the issues and important factors that need to be taken into considerations when designing, training, and validating new methods. We also point out some of the limitations of prior works, persistent challenges, and the aspects that require further investigation.
III-A Validation
@@
A shortcoming of many of the studies cited above is with regard to validation. We break down the main aspects of this limitation below.

### v2

@@ IV-J Segmentation
Tissue segmentation in the dMRI space is needed for anatomically constrained tractography and structural connectivity analysis [219, 213]. Given the lower spatial resolution and tissue contrast of dMRI, this segmentation is usually obtained via segmenting anatomical MRI data followed by registration to the dMRI space. This registration can introduce significant error because the contrast and image distortions between the modalities can be very different. Therefore, direct tissue segmentation in dMRI is desirable. Machine learning methods, and in particular deep learning techniques, have shown great promise here as well.
@@
Prior to the development of deep learning-based segmentation methods, classical machine learning methods such as fuzzy c-means clustering with spatial constraints [346] and sparse representation of the dMRI signal in dictionaries [347] were used to segment the brain tissue. Ciritsis et al. determined the optimal set of dMRI measurements and DTI features for tissue classification with support vector machines (SVM) [348]. Schnell et al. computed rotation-invariant features from HARDI measurements and used them as input to an SVM to segment the brain tissue into six classes. The class labels included white matter with parallel (single) fibers and white matter with crossing fibers [349].
@@
More recently, Golkov et al. [16] extended their deep learning tissue microstructure estimation technique to develop a segmentation method that mapped the dMRI signal in each voxel to a probability for white matter (WM), gray mater (GM), cerebrospinal fluid (CSF), and MS lesion. Their simple and flexible approach allowed them to include an extra FLAIR channel, which could be especially useful for MS lesion segmentation. Their model accurately segmented the brain tissue, and detected MS lesions with area under the receiver operator characteristic curve (AUC) in the range 0.88-0.94. The same authors also proposed a variational auto-encoder model for novelty detection based on q-space dMRI data and showed that it could accurately segment MS lesions with an AUC of approximately 0.90 [350]. Another study used DTI and DKI features as input to a multi-view CNN to achieve accurate segmentation of WM, GM, and CSF [351]. The model was trained on HCP-quality data, where the anatomical and dMRI data could be registered precisely. The trained model achieved high segmentation accuracy on clinical dMRI acquisitions. The same research team trained a deep learning model to parcellate the brain into 101 regions using DTI-derived parameter maps as input [352]. The authors showed the utility of the new method in improving tract-specific streamline bundle extraction from whole-brain tractograms. A recent study addressed the challenging task of fetal brain segmentation in dMRI [353]. This is a particularly difficult application due to the low contrast of fetal brain dMRI. The authors used DTI-derived parameters as the input to their model, which was based on vision transformers. They proposed a novel self-supervised learning approach that enabled them to train the model on a large unlabeled dataset and a smaller labeled dataset. This method achieved a mean DSC of 0.84-0.90 on different tissue types. Experiments showed that this method drastically improved fetal brain tractography. For isointense stage of brain development where white matter and gray matter have similar intensity in T1- and T2-weighted images (approximately between 6 and 8 months), one study used an additional FA channel to accurately segment the brain tissue with an FCN [354]. Segmentation of the intracranial volume also know as brain masking or skull stripping, which is typically performed as a pre-processing step, has also been addressed with machine learning methods [355, 356, 357, 358].
V Discussion
@@
This section presents some thoughts on the application of machine learning for dMRI analysis. It discusses important factors that need to be taken into considerations when designing, training, and validating new methods. Additionally, it points out some of the limitations of prior works, persistent challenges, and topics that require further investigation. Figure 2 shows the outline of the topics discussed in this section.
Figure 2: Outline of the main aspects of application of machine learning in dMRI that are discussed in Section V.
V-A Validation
@@
A shortcoming of many of the studies discussed above is with regard to validation. This subsection breaks down the main aspects of this limitation.
V-A1 Validation on abnormal/pathological data
The methods are usually trained and validated on data from normal brains. This can be a major flaw because brain pathologies may represent vastly different and heterogeneous structure and tissue microstructure that drastically alter the dMRI signal [211, 359, 360]. Brain pathologies such as tumors can also change the brain anatomy and impact the performance of automatic tract analysis methods [306, 305]. For tract delineation, one study has shown that on pathological brains conventional atlas-based methods work better than the more recent deep learning techniques [361]. Another study [362] found that the accuracy of automatic tract segmentation methods such as TractSeg dropped significantly (DSC=0.34) on brains with tumors. A new human-in-the-loop method based on active learning achieved a mean DSC of 0.71 on the same dataset.